OpenStudy (selena2345):

During one month in Maine, the highest temperature was 24 Celsius and the lowest temperature was -13 Celsius. What was the difference between the highest and lowest temperatures in Maine that month ? The next month the lowest temperature was -8 Celsius. What was the difference in the lowest temperatures between the two months ?

OpenStudy (anonymous):

@hartnn , this is to long. Take over

OpenStudy (selena2345):

Hartn take over

OpenStudy (anonymous):

thanks @hartnn

hartnn (hartnn):

just the question was long ... highest temperature was 24 Celsius the lowest temperature was -13 Celsius. difference = highest - lower = 24 - (-13) =...?

hartnn (hartnn):

hint : - * - =+ so, 24-(-13) = 24+13 =.... ?

OpenStudy (selena2345):

24-(-13)=37

hartnn (hartnn):

yes, thats the correct difference for first part :)

OpenStudy (selena2345):

okay

hartnn (hartnn):

for 2nd part, lowest temp this month = -13 lowest temperature next month = -8 difference = -8 -(-13) = ... ?

OpenStudy (selena2345):

-8-(-13)=5

hartnn (hartnn):

thats correct .

OpenStudy (selena2345):

okay